Grade 11Mechanics

A rocket of mass 40000kg is in empty space.Determine its velocity increase after burning all its fuel if the mass of the fuel is 90 % of the mass of the rocket.The speed of the exhaust gases relative to the rocket is 3000m/s

avanthika
8 Years agoGrade 11

1 Answer

Sandeep Bhamoo
ApprovedApproved Tutor Answer8 Years ago
Here mass of fuel burnt (mf)=(90/100)40000= 36000kg, mass of remaining rocket(mr)= (10/100)40000 =4000kg now by conservation of momentum ,36000(vf)= 4000(vr). Vr = 9Vf And since relative velocity =Vf+Vr = 10Vf=3000, Vf=300m/s, and Vr =2700m/s please do approve my answer if correct